OpenStudy (anonymous):

simplify your expression below m12-m6 =

OpenStudy (anonymous):

\[m^12-m^6=m^6(m^6-1)=m^6[(m^3)^2-1^2]=m^6[m^3)-1][m^3+1]\]

OpenStudy (anonymous):

Now \[[m^3)-1][m^3+1] can further be expanded by using (a^3+b^3) = (a+b)(a^2-ab+b^2) \] \[and (a^3-b^3) = (a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

so is it , A. m6 or B. m12 C. m-6 D. m18

OpenStudy (anonymous):

I am still confused..sorry

OpenStudy (anonymous):

\[m^6(m^6−1)\]

OpenStudy (anonymous):

thank you